OpenStudy (anonymous):

What is the solution to the equation log 2 x – log 2 4 = 2 ? x = 1 x = 18 x = 8 x = 16

OpenStudy (mertsj):

\[\log_{2}x-\log_{2}4=2 or \log_{2}\frac{x}{4}=2 \]

OpenStudy (mertsj):

Now put that in exponential form.

OpenStudy (anonymous):

SO 1 WOULD BE MY ANSWER?

OpenStudy (mertsj):

What is the exponential form of that equation? Hint: The base is 2 and the exponent is also 2

OpenStudy (anonymous):

IM STUCK

OpenStudy (mertsj):

\[2^2=\frac{x}{4}\]

OpenStudy (mertsj):

That is the exponential form. Solve it.

OpenStudy (anonymous):

X=8?

OpenStudy (ash2326):

\[2^2=\frac x 4\] \[4=\frac x 4\] Now solve for x

OpenStudy (anonymous):

8 RIGHT

OpenStudy (ash2326):

How do you get 8? \[\frac x 4= 4\] \[x=4\times 4\]

OpenStudy (anonymous):

16

OpenStudy (ash2326):

Yes, we can check if it's correct \[\log_2 x-\log_2 4=2\] \[\log_2 16 -\log_2 4 =2\] \[4-2=2\] So it's correct Did you understand ?

OpenStudy (anonymous):

KIND OF